### 27 Dec 2023 v1

# Computational Tradeoffs of Optimization-Based Bound Tightening in ReLU Networks

Abstract
The use of Mixed-Integer Linear Programming (MILP) models to represent neural networks with Rectified Linear Unit (ReLU) activations has become increasingly widespread in the last decade. This has enabled the use of MILP technology to test—or stress—their behavior, to adversarially improve their training, and to embed them in optimization models leveraging their predictive power. Many of these MILP models rely on activation bounds. That is, bounds on the input values of each neuron. In this work, we explore the tradeoff between the tightness of these bounds and the computational effort of solving the resulting MILP models. We provide guidelines for implementing these models based on the impact of network structure, regularization, and rounding.
Keywords:
Activation Bounds Neural Networks Neural Surrogate Models Mixed-Integer Linear Programming Rectified Linear Units.
1 Introduction
One of the key sources of the expressiveness of neural networks is their nonlinearity, which comes from the activation functions. The most commonly used activation function in the last decade is the Rectified Linear unit (ReLU) activation function [21, 34, 18, 28, 38], defined as $\sigma(x)=\max\{0,x\}$. All ReLU neural networks represent piecewise linear functions [36]. Conversely, all piecewise linear functions can be represented as ReLU networks [3]. Moreover, ReLU networks are universal function approximators under a variety of architectures [49, 31, 22, 35], meaning that they can represent almost any function to arbitrary precision in the same way as with other activation functions [11, 17, 23].
This piecewise linear structure of the ReLU activation function facilitates embedding these networks in mixed-integer linear models. For an input $x\in[\ell,u]$, we can represent such $x$$y$ relation $y=\sigma(x)$ as the following linear system.
$x=y-h,\,y\leq uz,\,h\leq-\ell(1-z),\,y\geq 0,h\geq 0,\,z\in\{0,1\}.$ (1)
Here, $z$ indicates if $x\geq 0$ or not. This system can model every neuron using ReLU, thus effectively representing the neural network evaluation as a system of linear inequalities over mixed-integer variables.
These models have been used in several applications involving trained neural networks in one way or another. The applications focused on the neural networks themselves include the generation of adversarial inputs [10, 16, 27], verification and robustness certification [44, 29, 30], feature visualization [16], counting linear regions [42, 40, 7], compression [41, 43, 14], providing strong convex relaxations [1, 2, 46], generating heuristic solutions for neural surrogate models [37, 45], producing counterfactual explanations [25], and constraining reinforcement learning policies [12, 5]. We refer the reader to the recent survey [24] for a review of further uses of polyhedral theory and MILP technology in the specific context of deep learning. Moreover, modeling frameworks [4, 8, 15, 32, 19] have been inspired by the wider variety of problems beyond deep learning, e.g., in power grids [9, 33], automated control [39, 47, 48], vehicle routing [12], among others.
An important aspect of these models is that they require valid bounds on the input of each ReLU function, as in (1). While the tightness of these bounds does not affect the correctness of the formulation, exceedingly large bounds may negatively affect the computational performance when solving the resulting MILP models. This is because the effectiveness of MILP technologies largely depends on the strength of LP relaxation bounds.
In this work, we study the tradeoffs between the time spent computing tight activation bounds and the time gained from using such bounds in mixed-integer linear models involving trained neural networks. First, we analyze how much more expensive it is to compute tight bounds using an Optimization-Based Bound Tightening (OBBT) approach compared to a simpler interval-based propagation approach. We evaluate runtimes according to different characteristics of the neural networks, such as architectural structure, regularization during training, and network pruning level. Second, we test these different bounds when applied to network verification, hence characterizing the aforementioned tradeoff. Our goal is to provide computational guidelines for researchers and practitioners when using MILP models involving trained neural networks.
Though many papers tend to use MILP optimization for bound tightening  [16, 40, 46, 6], to our knowledge there is very little work analyzing this computational tradeoff. An exception is the work of Liu et al. [30], which compares various algorithms for computing activation bounds, some of which we also consider here. In this paper, we complement the work of [30] in two important ways: (i) we include a sensitivity analysis of different bound computation approaches with respect to multiple network characteristics; and (ii) we analyze the effect of the bounds quality in other optimization problems.
2 Background
2.1 Neural Networks
A neural network is a function $f$ defined over a directed graph that maps inputs $x\in\mathbb{R}^{n_{0}}$ to $f(x)\in\mathbb{R}^{n_{L}}$. The directed graph $G=(V,E)$ representing the network decomposes into layers $V=\bigcup_{i\in[L]_{0}}V_{i}$; $V_{0}$ is the input layer and $V_{L}$ as the output layer. Layers $1$ through $L-1$ are the hidden layers. In this work, we consider fully connected feed-forward networks, meaning that all nodes in layer $V_{j}$ are connected to all nodes in layer $V_{j+1}$ and that arcs only connect nodes in adjacent layers. This is mainly to simplify the discussion, as this framework can also be adapted to other feed-forward architectures.
The neural network computes an output in the following way. Each layer $l$ outputs a vector $h^{(l)}$ to the following layer, starting with $h^{(0)}=x$. Then we apply an affine transformation ${W^{(l)}}^{\top}h^{(l)}+b^{(l)}=:a^{(l+1)}$. Finally, the activation function is applied (component-wise) to $a^{(l+1)}$ to produce the output of layer $l+1$, i.e., $h^{(l+1)}=\sigma(a^{(l+1)})$. The last layer often applies an extra transformation, such as $\mbox{softmax}(z)=e^{z_{i}}/\sum_{j=1}^{w}e^{z_{j}}$, to model a probability distribution.
2.2 Computing Activation Bounds
If the input $x\in\mathbb{X}$ is bounded, it is easy to see that each value of $a^{(l)}$ is bounded as well; these activation bounds are needed for a formulation using (1). Fix a layer $m$ and suppose we have computed values $\mbox{LB}_{j}^{(l)},\mbox{UB}_{j}^{(l)}$ such that $a_{j}^{(l)}\in[\mbox{LB}_{j}^{(l)},\mbox{UB}_{j}^{(l)}]$ for $l\leq m-1$. Then we can compute a valid bound for neuron $n$ as follows:
$\mbox{UB}_{n}^{(m)}=\max_{x\in\mathbb{X}}\quad$ $a^{(m)}_{n}$ (2a)
s.t. ${W^{(l)}}^{\top}h^{(l)}+b^{(l)}=a^{(l+1)}$ $\forall l\in[m-1]_{0}$ (2b)
$a^{(l)}=h^{(l)}-\overline{h}^{(l)}$ $\forall l\in[m-1]$ (2c)
$h_{j}^{(l)}\leq\mbox{UB}_{j}^{(l)}z_{j}^{(l)}$ $\forall l\in[m-1],\forall j\in[n_{l}]$ (2d)
$\overline{h}_{j}^{(l)}\leq-\mbox{LB}_{j}^{(l)}(1-z_{j}^{(l)})$ $\forall l\in[m-1],\forall j\in[n_{l}]$ (2e)
$h^{(l)},\overline{h}^{(l)}\geq 0,\,z^{(l)}\in\{0,1\}^{n_{l}}$ $\forall l\in[m-1]$ (2f)
$h_{j}^{(0)}\in[\mbox{LB}_{j}^{(0)},\mbox{UB}_{j}^{(0)}]$ $\forall j\in[n_{0}]$ (2g)
Here $n_{l}:=|V_{l}|$ is the number of neurons in layer $l$ (the width). Note also that $\mbox{LB}_{n}^{(m)}$ can be computed by minimizing instead of maximizing. Solving (2.2) to determine $\mbox{LB}_{j}^{(l)},\mbox{UB}_{j}^{(l)}$ is NP-hard [26], and so it is unlikely that an efficient algorithm exists for solving (2.2). However, solving the linear relaxation of (2.2) also provides valid, quickly computable, but weaker bounds. Henceforth, we refer to these weaker bounds as $\widehat{\mbox{LB}}_{j}^{(l)}$ and $\widehat{\mbox{UB}}_{j}^{(l)}$.
Note that even weaker bounds can be computed, without the need to solve any optimization problem. This follows from Proposition 1, where upper activation bounds for a node can be computed from those in the previous layer. Lower activation bounds can be computed by reversing the sign of the bounds. If we do this for all the layers iteratively, we will get valid bounds for the entire network.
Proposition 1 (Naive Bound)
[…]
To find all activation bounds using formulation (2.2), we proceed forward from the input layer to compute upper and lower activation bounds one layer at a time. This procedure is described in Algorithm 1. We say that the bounds $\mbox{LB}_{n}^{(m)}$ and $\mbox{UB}_{n}^{(m)}$ resulting from this algorithm are the strong node activation bounds. Note that in step 1 we can solve the relaxed version of (2.2) if we want to compute $\widehat{\mbox{LB}}_{j}^{(l)}$ and $\widehat{\mbox{UB}}_{j}^{(l)}$. We call these the weak node activation bounds. If, instead, we ignore step 1 altogether, we obtain what we call the naive node activation bounds.
We remark that in our implementation of Algorithm 1, we parallelized the for loop in line 5; this is possible because problem (2.2) is independent $\forall n\in[n_{m}]$ for a fixed layer $m\in[L]$. Note also that if the MILP solver fails to produce an optimal integer solution, we ask the solver to provide us with the best dual bound available, if there is one, before using the bound from Proposition 1.
Input: Trained neural network, input layer bounds $\mbox{LB}^{(0)},\mbox{UB}^{(0)}$
1 Get admissible bounds $\mbox{LB}^{(l)},\mbox{UB}^{(l)}$ $\forall l\in[L]$ using Proposition 1
2 Initialize the variables of (2.2) and add (2g)
3 for $1\leq m\leq L$ do
[…]
5       for $1\leq n\leq n_{m}$ do
6             Solve (2.2) and its minimization versions for the current node $(n,m)$
7             if Valid bounds for $a_{n}^{(m)}$ are found then
8                   Set $\mbox{UB}_{n}^{(m)}$ and $\mbox{LB}_{n}^{(m)}$ with the corresponding values
9
10            else
11                   Propagate worst bound of layer $m-1$ with Proposition 1.
12
13
Return $\left\{\mbox{LB}^{(l)},\mbox{UB}^{(l)}\right\}_{l=0}^{L}$
Algorithm 1 Bounder
2.3 Verification Problems
To analyze and compare the impact of the different ways of computing the activation bounds, we will use them in a verification problem formulation. In this problem, we start with reference input $x_{0}$ and formulate an optimization problem that finds a small perturbation that is classified differently to $x_{0}$.
For the sake of brevity, we will not display the full model, but it suffices to say that we simply modify the formulation (2.2) to have an objective function that maximizes the difference in the classification of $x_{0}$ and the perturbed input.
3 Experimental Setup
Our computational experiment consisted of three steps. First, we trained 96 neural networks using the MNIST dataset [13] as follows:
  1. 1.
    Architecture: We used four different neural network architectures, varying in the number of hidden layers $N_{l}$ and nodes per layer $N_{n}$, represented as $N_{l}\times N_{n}$; each layer being fully-connected and the last being a softmax layer.
  2. 2.
    Regularization and Training Duration: The networks were trained for 50 epochs with L1 and L2 regularization, employing a penalty parameter $\lambda$ set to values in $\{10^{-4},2\cdot 10^{-4},3\cdot 10^{-4}\}$.
  3. 3.
    Network Pruning: Post-training, networks were pruned using a threshold $\varepsilon$ within $\{10^{-4},10^{-3},10^{-2},10^{-1}\}$ to enhance numerical stability.
Second, for each node in each neural network we proceeded to compute strong and weak activation bounds using Algorithm 1 with Gurobi 9.1.2 [20] for MILP and LP runs. A time limit of one hour was used to compute each bound. We also compute the naive activation bounds.
Third, for each network architecture and each type of activation bounds (strong, weak, naive), we defined ten instances of the verification problem, defined in Section 2.3. For this, we select one network of each architecture with L1 regularization, a pruning threshold $\varepsilon=0.0002$, one random image of each class in MNIST as a reference $x_{0}$ and a radius $\epsilon=0.18$ to perturb the image. The resulting problem was then solved with Gurobi 9.1.2 [20]. A time limit of one hour was used to solve each problem. All runs were performed on Dell PowerEdge C6420 Intel Xeon Gold 6152 @ 2.10 GHz with 22 cores; we allocated 4 CPU cores and 6 GB of RAM per network.
4 Results
4.1 Strong Versus Weak Activation Bounds
Our first analysis seeks to determine how different the strong and weak bounds are to each other. This has also been reported in [30], but for the sake of completeness we also include our analysis. For this comparison, we use the relative optimality (RO) gap given by $\frac{|\widehat{LB}-LB|}{|LB|+10^{-10}}$, where $\widehat{LB}$ and $LB$ are the weak and strong lower activation bounds obtained by the LP and MIP, respectively.
In Figure 1, we display a summary of our results. Overall, we can see that the strong and weak bounds are remarkably similar to each other. Though not shown, results are similar for the upper activation bounds. In the worst case, we obtained a weak bound whose value was twice that of the corresponding strong bound (Figure 0(d)). We also observe an expected decay in the strength of the weak bound with respect to the depth of the network. We confirmed this with an analysis of the relative gap layer by layer, which we omit here due to space requirements. However, we remark that we obverse an exponential growth on RO gap by layer.
(a) 1-layer network.
(b) 2-layer network.
(c) 5-layer network.
(d) 10-layer network.
Figure 1: Relative optimality gap on the last layer versus rounding threshold $\varepsilon$ for each architecture and regularization type (L1/L2) and level $\lambda$.
(a) 1-layer network.
(b) 2-layer network.
(c) 5-layer network.
(d) 10-layer network.
Figure 2: Running times of Algorithm 1 for strong and weak bounds in the L1 case.
4.2 Solving Times
A second comparison we report is the running time (wall clock) on the previous experiments. In Figures 2 and 3, we show the total runtime for the L1 and L2 regularizations, respectively.
(a) 1-layer network.
(b) 2-layer network.
(c) 5-layer network.
(d) 10-layer network.
Figure 3: Running times of Algorithm 1 for strong and weak bounds in the L2 case.
In both regularization types, and for every regularization value $\lambda$, we can see the expected improvement in computation time for the weak bounds. This improvement is much more drastic on deeper networks as shown in Figures 1(c), 1(d), 2(c), 2(d). Also note that the pruning threshold has more impact on the strong bounds than on the weak ones, as shown in Figures 1(b), 1(c), 1(d), 2(b), 2(c), and 2(d). The same phenomenon happens with $\lambda$ in this case.
Overall, we observe that the weak bounds provide a good alternative for computing activation bounds. They can be computed efficiently, their running time scales well, and they have good quality with respect to the strong bounds. This was also noted in [30]. Besides confirming this, we also show that the weak bounds performance is robust across different values of pruning thresholds and regularization. In the following subsection, we test how a worse bound affects a verification model that depends on them.
4.3 Effectiveness On Verification Problems
To test the performance of different bound qualities, we consider three bound types to use in the verification problem from Section 2.3: the strong and weak bounds we discussed before, and naive bounds given by Proposition 1.
(a) Instances that reach optimality.
(b) Instances that reach a gap $\leq 1\%$
Figure 4: Comparison for the different bound types used.
In Figure 3(a), we see that strong bounds got the most optimal solutions, followed by weak bounds and naive bounds. This was expected. What is more interesting is that if we take a closer look at the instances that were near optimality (gap lower than $1\%$), as shown in Figure 3(b), we can see that weak bounds have a better performance overall, even surpassing strong bounds in some ranges.
Figure 5: Instances that reach a gap $1\%$ by network architecture.
In Figure 5, we disaggregate the result based on the network architecture. Note that all the bound types reach a gap lower than 1% for shallower networks. We also note that naive bounds work quite well for these networks. For networks such as 2x800 in Figure 5, we can see that weak bounds surpass strong ones, despite taking much lower time to compute. Finally, we see that strong bounds give the best results for the deepest networks overall.
5 Conclusions
In terms of the computation of activation bounds, we see that having some properties such as a regularized or pruned network will facilitate their computation. In addition, the weak bounds proved to be comparable to strong bounds, but with considerably less computing time and more stability across different architectures. Computing the bounds with LP relaxations provides a good compromise of quality in favor of computability.
On the performance of different bound types on a verification problem, the results show that using the weak bounds provides better results overall, and could be used as a general “rule of thumb”. However, here the dominance is not as clear across different contexts. For example, on shorter networks naive bounds work as well as the other bound types, and on deeper networks the performance of the weak bounds tends to get worse relative to strong bounds. These results could be used to design a better approach; for example, using a hybrid method that uses naive bounds on the first layers, then weak bounds on the middle layers and strong bounds for the deeper layers.
{credits}
5.0.1 Acknowledgements
G.M. was supported by ANID Fondecyt grant 1231522 and by ANID PIA/PUENTE AFB230002. M.G. was supported by ANID Fondecyt grant 1231092 and Centro de Modelamiento Matemático (CMM), FB210005, BASAL funds for centers of excellence from ANID-Chile. Powered@NLHPC: This research was partially supported by the supercomputing infrastructure of the NLHPC (CCSS210001). T.S. was supported by the National Science Foundation (NSF) grant IIS 2104583.